A father is three times his daughter's age. In 10 years time the sum of their ages will be 68. How old are the father and daughter?
Answer by Cromlix(4381)

Hi there,
A father is three times his daughter's age.
F = 3D
In 10 years time the sum of their ages will be 68.
F + 10 + D + 10 = 68
Replace F with 3D
3D + 10 + D + 10 = 68
Collect like terms
3D + D = 68 - 10 - 10
4D = 48
D = 12
Daughter = 12yrs old
Father = 36yrs old
Hope this helps you :-)